# Version dated: November 6, 2025

### v1

@@ 4.3 HMC with an approximate gradient
In the current work, we employ a Metropolis-within-Gibbs MCMC framework to sample from the full posterior distribution in Equation 24 (metropolis_equation_1953; geman_stochastic_1984). Within this framework, we partition the parameters into two blocks: the CTMC parameters, $\boldsymbol{\theta}$, and a second block containing all other parameters. While we adopt standard random-walk kernels to propose the second block of parameters following the rich and established literature in Bayesian phylogenetics (hassler_data_2023), we use an HMC sampler with surrogate trajectories (duane_hybrid_1987; li_neural_2019) for the CTMC parameters as detailed in Section 3.
@@
To make the HMC operational, we proceed by computing the explicit form for the derivative of the tree data likelihood function using the aforementioned approximation of the matrix exponential gradient. As we showed for the case of a sequentially observed CTMC with known sampling times (Section 3.1.2), also for a tree-based likelihood we can exploit the sparsity of the matrix $\mathbf{E}_{ij}$ in the gradient approximation from Equation 18 to further reduce the computational complexity and derive an easily operationalizable formula. Specifically, based on the calculations shown in Appendix 6.1, the gradient of the tree data likelihood with respect to the normalized rates can be written as follows (magee_random-effects_2024):
@@
$\frac{\partial p\left(\mathbf{y}\right)}{\partial\lambda_{ij}}$ $=\sum_{u=1}^{2N-2}\mathbf{p}_{u}^{\prime}\left[\frac{\partial e^{t_{u}\mathbf{\Lambda}}}{\partial\lambda_{ij}}\right]^{\prime}\tilde{\mathbf{q}}_{\text{pa}(u)}+\mathbf{p}_{2N-1}^{\prime}\frac{\partial}{\partial\lambda_{ij}}\boldsymbol{\pi}_{\text{root}},$ (28)
@@
where $\tilde{\mathbf{q}}_{\text{pa}(u)}:=\left[\mathbf{q}_{\text{pa}(u)}\circ{\mathbf{P}}\hskip-1.99997pt\left(t_{\text{sibl}(u)}\right)\mathbf{p}_{\text{sibl}(u)}\right]$. Substituting $\frac{\partial e^{t_{u}\mathbf{\Lambda}}}{\partial\lambda_{ij}}$ by the approximation presented in Equation 18, we can re-write the first summation on the right hand side of Equation 28 in a substantially simpler form:
@@
$\sum_{u=1}^{2N-2}\mathbf{p}_{u}^{\prime}\left[\frac{\partial e^{t_{u}\mathbf{\Lambda}}}{\partial\lambda_{ij}}\right]^{\prime}\tilde{\mathbf{q}}_{\text{pa}(u)}$ $=\sum_{u=1}^{2N-2}\mathbf{p}_{u}^{\prime}\left[t_{u}e^{t_{u}\mathbf{\Lambda}}\mathbf{E}_{ij}\right]^{\prime}\tilde{\mathbf{q}}_{\text{pa}(u)}$
$=\sum_{u=1}^{2N-2}t_{u}\mathbf{p}_{u}^{\prime}\mathbf{E}_{ij}^{\prime}\left[e^{t_{u}\mathbf{\Lambda}}\right]^{\prime}\tilde{\mathbf{q}}_{\text{pa}(u)}$
$=\sum_{u=1}^{2N-2}t_{u}\mathbf{p}_{u}^{\prime}\mathbf{E}_{ij}^{\prime}\mathbf{q}_{u}$
$=\sum_{u=1}^{2N-2}t_{u}p_{uj}q_{ui},$ (29)
@@
where $p_{uj}$ and $q_{ui}$ are the components respectively of $\mathbf{p}_{u}$ and $\mathbf{q}_{u}$, and we apply Equation 27 in the third line. The approximation reduces the computational complexity for computing $\frac{\partial p\left(\mathbf{y}\right)}{\partial\lambda_{ij}}$ from ${\cal O}(NK^{3})$ using the traditional matrix exponential gradient computation (kalbfleisch_analysis_1985) to just ${\cal O}(N)$, where $N$ stems from the summation in Equation 4.3. Since we have to repeat these calculations for each entry of the infinitesimal matrix, the total computational complexity of our proposed method becomes of order ${\cal O}(NK^{2})$ versus ${\cal O}(NK^{5})$ for the exact method.
@@ 4.4.1 Simulation
To assess the performance of the proposed methods for parameter estimation, we conduct a simulation study. Based on faria_toward_2011, we begin by considering a CTMC describing the transmission process of the rabies virus infecting $17$ host bat species. Each pair of species is associated with their genetic distances.
For simulation purposes, we set the log-rates of the CTMC equal to a quadratic function of those genetic distances as shown in Figure 1 (dashed line). Next, we select a phylogenetic tree sampled from the posterior distribution given aligned genomic sequences obtained in faria_toward_2011, containing $372$ tips and spanning an evolutionary history of approximately 270 years. Subsequently, we simulate a single alignment (one observation per tip) by evolving the CTMC along the tree branches, starting from the root. The evolutionary rate, representing the average “speed” of evolution, is fixed at $2\%$ per year. This corresponds to an expected total of roughly five transitions along the path from the root to the most recent tip.
To analyze the simulated data, we employ a GP prior on the log-rates, using a squared exponential kernel with the host-species genetic distances as a covariate. We place independent exponential priors with rate $1$ to the kernel hyperparameters (the marginal scale and the length-scale). As shown in Figure 1 (red line), the method recovers the underlying nonlinear log-rate function. As a comparison, we also run the analysis using a log-linear model (LL, blue line). Quantitatively, across posterior draws the GP achieves a median root mean square error (RMSE) of $0.170$ ($95\%$ highest posterior density interval (HPDI) 0.0390.3110.039–0.311), versus $0.487$ ($95\%$ HPDI 0.4040.5890.404–0.589) for the LL model—a $\approx 65\%$ reduction (LL $\approx 2.9\times$ larger), indicating superior accuracy. Moreover, pointwise $95\%$ HPDI coverage of the true log-rate function is $100\%$ for the GP and $28.68\%$ for the log-linear model, with average HPDI widths of $0.63$ and $0.39$, respectively; thus, the GP bands are slightly wider, as expected, but substantially better calibrated.
@@
Figure 1: Simulation Study
Note: The figure shows a performance comparison between the log-linear (LL) and Gaussian process (GP) models in recovering the CTMC log-rates in a simulated dataset. The dashed line represents the true log-rates, which were set as a quadratic function of the host genetic distances. The solid lines show the posterior median normalized log-rates inferred under the LL (blue) and the GP (red) models, while the shaded areas cover the $95\%$ highest posterior density intervals (HPDIs). Only slope uncertainty is depicted, as all values are normalized.
Figure 2: Computational Speed Comparison
Note: The figure shows a comparison of the average time (milliseconds, ms) per gradient evaluation as a function of the number of CTMC states between the approximate (red circles) and the exact numerical gradient (blue triangles). Dotted lines indicate theoretical scaling trends, ${\cal O}(K^{2})$ for the approximate and ${\cal O}(K^{5})$ for the numerical gradients, anchored at the estimated times for $128$ states.
@@
$q_{ij}=\operatorname{log}(x_{ij}+\operatorname{max}_{kl}(|x_{kl}|)+1),$ (30)
where the additive constants ensure non-negativity. This implies that the closer two states ($i,j)$ are, the higher their associated log-rates. Using this CTMC, we simulate one alignment over the same tree used in the previous simulation. For a logarithmically increasing sequence of states, $K=4,8,16,32,64,128$, we record the average computational time required per gradient evaluation. Figure 2 illustrates that the empirically observed times are consistent with the corresponding theoretical expectations for time complexity. In particular, we construct a log-log plot that juxtaposes the two empirical curves—one for the approximate gradient (red) and one for the numerical gradient (blue)—alongside two reference lines with slopes $2$ and $5$, respectively, which represent the theoretical scaling rates predicted for each method. The close alignment between the empirical trends and the reference slopes further corroborates the expected computational scaling behavior of the two gradient formulations.
4.4.2 Bat rabies viruses in North America
After validating our method on simulated data, we apply it to two real-world examples. We start by reproducing the analysis performed in faria_simultaneously_2013 of a dataset composed of $372$ rabies virus samples collected from $17$ species of infected bats in North America between 1997 and 2006 (streicker_host_2010). This study reconstructs the evolutionary history of the rabies viruses based on the sequences of the nucleoprotein gene, the location of the samples, as well as the species each host bat belongs to. Moreover, the authors investigate the transmission dynamics between different bat species in order to identify potential facilitating factors. Using a log-linear model, they conclude that cross-species genetic distance influences transmission: the more two bat species are genetically different, the less likely a transmission event.
We revisit these data, relaxing the log-linear assumption. Specifically, we assign to the infinitesimal log-rates a GP prior with a squared exponential kernel using the host genetic distance as a covariate. Moreover, we specify two exponential priors with rates $1$ and $2$, respectively, for the kernel scale and length hyperparameters. To avoid overfitting, we also constrain the length parameter to be at least one. As Figure 3(a) highlights, our results support the hypothesis that the genetic distance has a log-linear effect on the infinitesimal rates. The relation between this covariate and the rabies virus evolutionary history can also be visualized from a comparison of the two trees shown in Figure 4. On the left, the reader can see a viral tree that is produced according to the maximum credibility clade (MCC) criterion (Baele_HIPSTR_2025). This method selects the single tree that best represents the most well-supported clades from the full distribution of sampled trees, where a clade can be defined as a subset of the complete tree including a common ancestor and all of its descendants. Branches are colored based on the bat species most likely to harbor the virus throughout their corresponding period. On the right, a dendrogram illustrates the hierarchical clustering of bat species based on cross-species genetic distances, with groupings determined by Ward’s method (ward_jr_hierarchical_1963), that is, by minimizing the within-cluster variance. It is apparent that the two trees have a symmetric color-induced structure, suggesting a tendency for infection events to occur more frequently across more closely related species.
@@ 4.4.2 Bat rabies viruses in North America
Figure 3: Data examples: log-rates vs predictors
Note: These images illustrate the relationships between the log-rates and the predictors in two data sets. In particular, Figure 3(a) portrays the effect of the cross-species genetic distances in the rabies virus dataset, while Figure 3(b) focuses on the effect of the population density of the origin country in the transmission of the global flu. The solid lines represent the posterior median inferred log-rates under a log-linear (LL) and a Gaussian process-based (GP) models, while the shaded areas show the $95\%$ HPDIs. Notice that only the uncertainty in the slope is depicted, since all values are normalized.
Figure 4: Bat Rabies Data Example
Note: On the left, the viral tree shows the evolution of the rabies virus in a population of bats in North America. The yellow rectangular bars represent the $95\%$ HPDIs for the age of the corresponding node. The colors along each branch identify the most probable bat species hosting the rabies virus at that point in time. Besides, the tree on the right orders the bat species based on their genetic distance as outputted by Ward’s hierarchical clustering algorithm (ward_jr_hierarchical_1963). Comparing the two trees it stands out the symmetry in the distribution of the colors (bats species), supporting the hypothesis that the rabies viruses tend to be transmitted more across genetically more similar bat species.
@@ 4.4.3 Global Influenza and human transportation
As a second example, we use data on the global spread of seasonal H3N2 influenza between 2002 to 2007. The data consists of hemagglutinin sequences and sample locations from 1,441 human-sampled viral isolates, and were originally analyzed by lemey_unifying_2014 to study the annual spread of the virus across 14 global air communities. The authors’ original work used a log-linear model to guide the spatial dispersion process, leveraging covariates such as airplane seat numbers and origin country population densities.
@@
Figure 3(b) compares the inferred log-rates from both the log-linear and the GP models. The results suggest that the log-linear assumption is too restrictive, as it appears to overestimate the effect of low population densities while underestimating the effect of higher ones. Besides, Figure 5 shows the MCC tree reconstructing the evolutionary history of H3N2 Influenza, where each branch is colored with the air community where the corresponding lineage has spent more time. A quick inspection of the tree suggests that China and, in general, countries in Southern Asia were potential recurrent sources of infections. Indeed, their corresponding colors dominate the tree trunk, which we can here define as the sequence of external branches facing the top-right of the tree plot. To test this hypothesis, we compute the reward associated with each air community along the tree trunk; that is, the total expected time the CTMC spends in each of the states along the tree trunk (minin_counting_2007). The dot plot at the bottom of Figure 5 shows the posterior median trunk rewards associated with each air community using the GP model. In particular, the figure shows that the highest median trunk reward is associated with China (around 4 years between 2000 and 2007), confirming the aforementioned hypothesis.
Figure 5: Global Flu Data Example
Note: Maximum clade credibility (MCC) tree showing the evolutionary history of H3N2 Influenza. The yellow bars placed on the internal nodes represent the $95\%$ HPDIs for the node age. The different colors represent distinct air communities as represented on the map on the top-right, where the dots correspond to the analysed airports. Each branch color refers to the air community where the GP model predicts most time is spent. The dot plot with error bars on the bottom shows the trunk rewards in years based on the GP model. The trunk rewards accumulate the amount of time (in years) the CTMC spends in each state along the trunk, which in turn is defined as the set of external branches that face the top-right of the tree plot. The dot plot error bars show the $95\%$ HPDIs.

### v2

@@ 4.3 HMC with an approximate gradient
In the current work, we employ a Metropolis-within-Gibbs MCMC framework to sample from the full posterior distribution in Equation 17 (Metropolis et al., 1953; Geman and Geman, 1984). Within this framework, we partition the parameters into two blocks: the CTMC parameters, $\boldsymbol{\theta}$, and a second block containing all other parameters. While we adopt standard random-walk kernels to propose the second block of parameters following the rich and established literature in Bayesian phylogenetics (Hassler et al., 2023), we use an HMC sampler with surrogate trajectories (Duane et al., 1987; Li et al., 2019) for the CTMC parameters as detailed in Section 3.
@@
To make the HMC operational, we proceed by computing the explicit form for the derivative of the tree data likelihood function using the aforementioned approximation of the matrix exponential gradient. As we showed for the case of a sequentially observed CTMC with known sampling times (Section 3.1.2), also for a tree-based likelihood we can exploit the sparsity of the matrix $\mathbf{E}_{ij}$ in the gradient approximation from Equation 14 to further reduce the computational complexity and derive an easily operationalizable formula. Specifically, based on the calculations shown in Section S3 of the Supplement, the gradient of the tree data likelihood $p\left(\mathbf{y}\right)$ with respect to the normalized rates can be written as follows (Magee et al., 2024):
@@
$\frac{\partial p\left(\mathbf{y}\right)}{\partial\lambda_{ij}}$ $=\sum_{u=1}^{2N-2}\mathbf{p}_{u}^{\prime}\left[\frac{\partial e^{b_{u}\mathbf{\Lambda}}}{\partial\lambda_{ij}}\right]^{\prime}\tilde{\mathbf{q}}_{\text{pa}(u)}+\mathbf{p}_{2N-1}^{\prime}\frac{\partial}{\partial\lambda_{ij}}\boldsymbol{\pi}_{\text{root}},$ (21)
@@
where $\tilde{\mathbf{q}}_{\text{pa}(u)}:=\left[\mathbf{q}_{\text{pa}(u)}\circ{\mathbf{P}}\hskip-1.99997pt\left(b_{\text{sibl}(u)}\right)\mathbf{p}_{\text{sibl}(u)}\right]$. Substituting $\frac{\partial e^{b_{u}\mathbf{\Lambda}}}{\partial\lambda_{ij}}$ by the approximation presented in Equation 14, we can re-write the first term on the right hand side in a substantially simpler form:
@@
$\sum_{u=1}^{2N-2}\mathbf{p}_{u}^{\prime}\left[\frac{\partial e^{b_{u}\mathbf{\Lambda}}}{\partial\lambda_{ij}}\right]^{\prime}\tilde{\mathbf{q}}_{\text{pa}(u)}$ $=\sum_{u=1}^{2N-2}\mathbf{p}_{u}^{\prime}\left[b_{u}e^{b_{u}\mathbf{\Lambda}}\mathbf{E}_{ij}\right]^{\prime}\tilde{\mathbf{q}}_{\text{pa}(u)}$
$=\sum_{u=1}^{2N-2}b_{u}\mathbf{p}_{u}^{\prime}\mathbf{E}_{ij}^{\prime}\left[e^{b_{u}\mathbf{\Lambda}}\right]^{\prime}\tilde{\mathbf{q}}_{\text{pa}(u)}$
$=\sum_{u=1}^{2N-2}b_{u}\mathbf{p}_{u}^{\prime}\mathbf{E}_{ij}^{\prime}\mathbf{q}_{u}$
$=\sum_{u=1}^{2N-2}b_{u}p_{uj}q_{ui},$ (22)
@@
where $p_{uj}$ and $q_{ui}$ are the components respectively of $\mathbf{p}_{u}$ and $\mathbf{q}_{u}$, and we apply Equation 20 in the third line. The approximation reduces the computational complexity for computing $\frac{\partial p\left(\mathbf{y}\right)}{\partial\lambda_{ij}}$ from ${\cal O}(NK^{3})$ using the traditional matrix exponential gradient computation (Kalbfleisch and Lawless, 1985) to just ${\cal O}(N)$, where $N$ stems from the summation in Equation 4.3. Since we have to repeat these calculations for each entry of the infinitesimal matrix, the total computational complexity of our proposed method becomes of order ${\cal O}(NK^{2})$ versus ${\cal O}(NK^{5})$ for the exact method.
@@ 4.4.1 Simulation
To assess the performance of the proposed methods for parameter estimation, we conduct a simulation study. Based on Faria et al. (2011), we begin by considering a CTMC describing the transmission process of the rabies virus infecting $17$ host bat species. Each pair of species is associated with their genetic distances.
For simulation purposes, we set the log-rates of the CTMC equal to a quadratic function of those genetic distances as shown in Figure 1(a) (dashed line). Next, we select a phylogenetic tree sampled from the posterior distribution given aligned genomic sequences obtained in Faria et al. (2011), containing $372$ tips and spanning an evolutionary history of approximately 270 years. Subsequently, we simulate a single alignment (one observation per tip) by evolving the CTMC along the tree branches, starting from the root. The evolutionary rate, representing the average “speed” of evolution, is fixed at $2\%$ per year. This corresponds to an expected total of roughly five transitions along the path from the root to the most recent tip.
To analyze the simulated data, we employ a GP prior on the log-rates, using a squared exponential kernel with the host-species genetic distances as a covariate. We place independent exponential priors with rate $1$ to the kernel hyperparameters (the marginal scale and the length-scale). As shown in Figure 1(a) (red line), the method recovers the underlying nonlinear log-rate function. As a comparison, we also run the analysis using a log-linear model (LL, blue line). Quantitatively, across posterior draws the GP achieves a median root mean square error (RMSE) of $0.170$ ($95\%$ highest posterior density interval (HPDI) 0.0390.3110.039–0.311), versus $0.487$ ($95\%$ HPDI 0.4040.5890.404–0.589) for the LL model—a $\approx 65\%$ reduction (LL $\approx 2.9\times$ larger), indicating superior accuracy. Moreover, pointwise $95\%$ HPDI coverage of the true log-rate function is $100\%$ for the GP and $28.68\%$ for the log-linear model, with average HPDI widths of $0.63$ and $0.39$, respectively; thus, the GP bands are slightly wider, as expected, but substantially better calibrated.
(a) Statistical recovery of log-rates.
(b) Computational speed comparison.
Figure 1: Simulation studies. Panel (a) compares log-linear ($L$) and Gaussian process (GP) models in recovering CTMC log-rates from simulated data where true log-rates (dashed line) follow a quadratic function of host genetic distance. Solid lines show posterior median normalized log-rates; shaded regions represent 95% highest posterior density intervals (HPDIs). Only slope uncertainty is depicted due to normalization. Panel (b) compares computational efficiency, showing average time per gradient evaluation versus the number of CTMC states. The approximate gradient (red) scales as ${\cal O}(K^{2})$ while the exact numerical gradient (blue) scales as ${\cal O}(K^{5})$, with scaling trends anchored at 128 states.
@@
$q_{ij}=\log(x_{ij}+\max_{kl}(|x_{kl}|)+1),$ (23)
where the additive constants ensure non-negativity. This implies that the closer two states ($i,j)$ are, the higher their associated log-rates. Using this CTMC, we simulate one alignment over the same tree used in the previous simulation. For a logarithmically increasing sequence of states, $K=4,8,16,32,64,128$, we record the average computational time required per gradient evaluation. Figure 1(b) illustrates that the empirically observed times are consistent with the corresponding theoretical expectations for time complexity. In particular, we construct a log-log plot that juxtaposes the two empirical curves—one for the approximate gradient (red) and one for the numerical gradient (blue)—alongside two reference lines with slopes $2$ and $5$, respectively, which represent the theoretical scaling rates predicted for each method. The close alignment between the empirical trends and the reference slopes further corroborates the expected computational scaling behavior of the two gradient formulations.
@@
Figure 2: Data examples: log-rates vs predictors. Panel (a) shows the effect of cross-species genetic distances on rabies virus transmission rates; panel (b) shows the effect of origin country population density on global flu transmission. Solid lines represent posterior median inferred log-rates under log-linear ($L$) and Gaussian process-based (GP) models; shaded areas show 95% HPDIs. Only slope uncertainty is depicted due to normalization.
4.4.2 Bat rabies viruses in North America
After validating our method on simulated data, we apply it to two real-world examples. We start by reproducing the analysis performed in Faria et al. (2013) of a dataset composed of $372$ rabies virus samples collected from $17$ species of infected bats in North America between 1997 and 2006 (Streicker et al., 2010). This study reconstructs the evolutionary history of the rabies viruses based on the sequences of the nucleoprotein gene, the location of the samples, as well as the species each host bat belongs to. Moreover, the authors investigate the transmission dynamics between different bat species in order to identify potential facilitating factors. Using a log-linear model, they conclude that cross-species genetic distance influences transmission: the more two bat species are genetically different, the less likely a transmission event.
We revisit these data, relaxing the log-linear assumption. Specifically, we assign to the infinitesimal log-rates a GP prior with a squared exponential kernel using the host genetic distance as a covariate. Moreover, we specify two exponential priors with rates $1$ and $2$, respectively, for the kernel scale and length hyperparameters. To avoid overfitting, we also constrain the length parameter to be at least one. As Figure 2(a) highlights, our results support the hypothesis that the genetic distance has a log-linear effect on the infinitesimal rates. The relation between this covariate and the rabies virus evolutionary history can also be visualized from a comparison of the two trees shown in Figure 3. On the left, the reader can see a viral tree that is produced according to the maximum credibility clade (MCC) criterion (Baele et al., 2025a).
[…]
Figure 3: Bat Rabies data example. The viral tree (left) shows the rabies virus evolution in North American bats. Yellow bars represent 95% HPDIs for node ages; branch colors identify the most probable host bat species. The bats species tree (right) orders bat species by genetic distance using Ward’s hierarchical clustering algorithm (Ward Jr., 1963). Color symmetry between trees suggests rabies transmission occurs preferentially among genetically similar bat species.
This method selects the single tree that best represents the most well-supported clades from the full distribution of sampled trees, where a clade can be defined as a subset of the complete tree including a common ancestor and all of its descendants. Branches are colored based on the bat species most likely to harbor the virus throughout their corresponding period. On the right, a dendrogram illustrates the hierarchical clustering of bat species based on cross-species genetic distances, with groupings determined by Ward’s method (Ward Jr., 1963), that is, by minimizing the within-cluster variance. It is apparent that the two trees have a symmetric color-induced structure, suggesting a tendency for infection events to occur more frequently across more closely related species.
@@ 4.4.3 Global Influenza and human transportation
As a second example, we use data on the global spread of seasonal H3N2 influenza between 2002 to 2007. The data consists of hemagglutinin sequences and sample locations from 1,441 human-sampled viral isolates, and were originally analyzed by Lemey et al. (2014) to study the annual spread of the virus across 14 global air communities. The authors’ original work used a log-linear model to guide the spatial dispersion process, leveraging covariates such as airplane seat numbers and origin country population densities.
@@
Figure 2(b) compares the inferred log-rates from both the log-linear and the GP models. The results suggest that the log-linear assumption is too restrictive, as it appears to overestimate the effect of low population densities while underestimating the effect of higher ones.
[…]
Figure 4: Global Flu data example. Maximum clade credibility (MCC) tree showing H3N2 influenza evolutionary history. Yellow bars on internal nodes represent 95% HPDIs for node ages. Branch colors indicate distinct air communities (map, top-right; dots represent analyzed airports), with each color showing the air community where the GP model predicts most time is spent. The bottom plot shows trunk rewards (cumulative time in years the CTMC spends in each state along the trunk, defined as external branches facing top-right) based on the GP model, with error bars representing 95% HPDIs.
Besides, Figure 4 shows the MCC tree reconstructing the evolutionary history of H3N2 Influenza, where each branch is colored with the air community where the corresponding lineage has spent more time. A quick inspection of the tree suggests that China and, in general, countries in Southern Asia were potential recurrent sources of infections. Indeed, their corresponding colors dominate the tree trunk, which we can here define as the sequence of external branches facing the top-right of the tree plot. To test this hypothesis, we compute the reward associated with each air community along the tree trunk; that is, the total expected time the CTMC spends in each of the states along the tree trunk (Minin and Suchard, 2007). The dot plot at the bottom of Figure 4 shows the posterior median trunk rewards associated with each air community using the GP model. In particular, the figure shows that the highest median trunk reward is associated with China (around 4 years between 2000 and 2007), confirming the aforementioned hypothesis.